Allison drives 30 miles per hour through the city and 55 miles per hour on the rt. 95. she drove 90 miles from Boston to the York Beaches. How much of the time was city driving if she spent 1.2 hours on rt. 95?

Answer by stanbon(75887)

Allison drives 30 miles per hour through the city and 55 miles per hour on the rt. 95. she drove 90 miles from Boston to the York Beaches. How much of the time was city driving if she spent 1.2 hours on rt. 95?
---------------------------
Rt. 95 driving DATA:
Rate = 55 mph ; time = 1.2 hrs ; distance = rt = 66 miles
--------------------
City driving DATA:
Rate = 30 mph ; distance = 90-66 = 24 miles ; time = d/r = 24/30 = 4/5 hr
--------------------
Answer: 4/5 hr = 48 minutes
==============
Cheers,
Stan H.